# Meromorphic functions whose action on their Julia sets is Non-Ergodic

### v2

@@ 3. Disjoint wandering sets of positive measure
Because $\Psi_{ij}$ is univalent on $G(D_{k})$, the radius of $G(D_{k})$ is less than $\pi/2$; that is,
@@
For each $k>0$, let $\mathcal{G}^{i}_{k}\subset\mathcal{G}^{i}$ be the collection of $S_{m,n}^{i}\subset Hor_{k}^{i}$. It will be convenient to use the index $\delta^{i}$ where $S_{\delta^{i}}=S_{m,n}^{i}$. Note that by definition, each such square lies between the vertical lines
$V_{n}=\{{\mathcal{Z}}=x^{i}_{N}+n\pi+iy,y>\alpha_{0}\}\hbox{ and }V_{n+1}=\{{\mathcal{Z}}=x^{i}_{N}+(n+1)\pi+iy),y>\alpha_{0}\}.$
@@
By lemmas 2.6 and 2.7, for each $j=1,\ldots,N$, $\Psi_{i,j}({\mathcal{R}}_{\delta^{i}}^{j})$ is a topological quadrilateral contained in the horizontal strip $Hor_{k+1}^{j}$. Let ${\mathcal{U}}^{j}(S_{\delta^{i}})$ denote the union of all the squares $S_{\delta^{j}}$ entirely contained in the interior of $\Psi_{i,j}({\mathcal{R}}_{\delta^{i}}^{j})$:
@@
Note that because all the squares have side length $\pi$,
@@
For each $j=1,\ldots,N$, let ${\mathcal{P}}_{\delta^{i}}^{j}={\mathcal{P}}^{j}_{\delta^{i}}(S_{\delta^{i}})=\Psi_{i,j}^{-1}({\mathcal{U}}^{j}(S_{\delta_{0}^{j}}))$. Since it is the pullback of a union of squares, each ${\mathcal{P}}^{j}_{\delta^{i}}$ is a topological quadrilateral in ${\mathcal{R}}_{\delta^{i}}^{j}$. Proposition 3.1 shows that the expansion factor for $\Psi_{i,j}$ on $S_{\delta^{i}}$ is at least $\alpha_{k}/(4\pi)$; this says that
@@
Denote the family of indices $({\boldsymbol{\iota}},{\boldsymbol{\sigma}})$ for which ${\mathcal{P}}_{\boldsymbol{\iota}}^{\boldsymbol{\sigma}}$ has a nonempty interior, and is contained in $Hor_{k}^{i_{0}}$, by $\mathcal{I}_{k}^{l}$ For readability, identify the index $(\boldsymbol{\iota},{\boldsymbol{\sigma}})\in\mathcal{I}_{k}^{l}$ with the set of points in ${\mathcal{P}}_{\boldsymbol{\iota}}^{\boldsymbol{\sigma}}$.
@@
(11) $Y^{l}=\cup_{(\boldsymbol{\iota},\boldsymbol{\sigma})\in\mathcal{I}^{l}}{\mathcal{P}}_{\boldsymbol{\iota}}^{\boldsymbol{\sigma}}\text{ and let }Y^{\infty}=\cap_{l=0}^{\infty}Y^{l}.$
@@
By definition, if $K$ is the set defined by $\mathcal{I}^{l}_{k}$, then $\Psi^{l+1}(K)$ is a union of squares in $Hor_{k+1+l}^{j_{l}}$. Denote this union by $W=\Psi^{l+1}(K)$. By definition,
$\Psi^{l+1}(W\cap Y^{l+1})=\cup_{i=1}^{N}\cup{\mathcal{P}}_{S_{\delta^{j_{l}}}}^{i}.$
By our choice of $k$, the distance of points in ${\mathcal{Z}}_{j_{1}}^{-1}(Hor_{k+1+l}^{j_{l}})$ to the orbits of the $\lambda_{i}$ is considerably larger than $2\pi$ so that branches of $\Psi^{-(l+1)}$ are well defined. Denote the one that maps $W$ back to $K$ by $\Xi$ and apply the Koebe distortion theorem. It says that for $\zeta,\xi\in W$, there exists a constant $C_{0}$ independent of $\Xi$, such that
@@
$\frac{m(W\setminus Y^{l+1})}{m(W)}=\frac{m(\Xi(\cup(S_{\delta^{j_{l}}}\setminus\cup_{i=1}^{l}{\mathcal{P}}_{\delta_{j_{l}}}^{i})))}{m(\Xi(\cup S_{\delta^{j_{l}}}))}\leq C_{0}^{2}\frac{\sum m(S_{\delta_{j_{l}}}\setminus\cup_{i=1}^{N}{\mathcal{P}}_{\delta^{j_{l}}}^{i})}{\sum m((S_{\delta^{j_{1}}})}\leq\frac{C}{\alpha_{|k|+l+1}},$
for some constant $C=(4\sqrt{2}\pi^{2}+6)C_{0}^{2}>0$. ∎
@@
    Every pair of points $({\mathcal{Z}},{\mathcal{Z}}^{\prime})\in(S_{\delta^{i_{0}}},S_{\delta^{i_{0}}}^{\prime}),$ satisfies $\Im|{\mathcal{Z}}|-|\Im{\mathcal{Z}}^{\prime}|\geq 2\alpha_{k-1}$.
@@
    $\Psi^{l}(W_{1})\in Hor^{i_{1}}_{k+l}$, $\Psi^{l}(W_{2})\subset Hor^{i_{2}}_{k+l}$ for some $i_{1},i_{2}\in\{1,\ldots,N\}$;
@@
The first assertion shows that both $W_{1}$ and $W_{2}$ are wandering sets and the second shows that their forward orbits are disjoint. Note that ${\mathcal{Z}}_{i_{0}}$ is analytic, so both ${\mathcal{Z}}_{i_{0}}^{-1}(W_{1}),{\mathcal{Z}}_{i_{0}}^{-1}(W_{2})$ have positive measure.
@@ 4. Main Theorems 2 and 3
Let $P_{f}=\cup_{i=1}^{N}\{\lambda_{i},f(\lambda_{i}),\ldots,f^{p_{i}-1}(\lambda_{i}),\infty\}$ be the post-singular set. As we proved above, $f$ is not ergodic. From the extended dichotomy of Bock [Bock], we know that for almost every point $z\in{\mathbb{C}}$, $\lim_{n\to\infty}d(f^{n}(z),P_{f})\to 0$. This implies that the $\omega$-limit set $\omega(z)\subset P_{f}$ and in terms of the auxiliary variables, it says that for each such $z$, there exists a fixed $i\in\{1,\ldots,N\}$, a sequence $n_{k}\to\infty$, and an $n_{k_{0}}$ such that for all $n_{k}\geq n_{k_{0}}$, $f^{n_{k}}(z)$ is in the asymptotic tract $T_{i}$ and $\Im{\mathcal{Z}}_{i}(f^{n_{k}}(z))\to\infty$.
The set $Y^{\infty}$ was defined in (11) as a subset of $\cup_{i=1}^{N}{\mathcal{Z}}_{i}$ and we showed in lemma 3.3 that it has positive measure. In order to prove Main Theorem 2, which is a result about points in ${\mathbb{C}}$, we need to pull $Y^{\infty}$ back to ${\mathbb{C}}$. To do this, define $E^{\infty}=\cup_{i=1}^{N}{\mathcal{Z}}_{i}^{-1}(Y^{\infty}\cap{\mathcal{Z}}_{i})$ and set $E=\cup_{n\in{\mathbb{N}}}f^{-n}(E^{\infty})$.
@@
The set $E$ in ${\mathbb{C}}$ has full measure; that is, $m({\mathbb{C}}\setminus E)=0$.
@@
Note that these inequalities define a set of strips $\widetilde{Hor^{i}_{l_{k}}}$ that is different from, but overlaps the strips $Hor^{i}_{n_{k}}$.
@@
Because the diameter of $\widetilde{Q}^{n_{k}}$ is less than $16\sqrt{2}\alpha_{k-1}$, the diameter of $F_{n_{k}}(\widetilde{Q}^{n_{k}})$ tends to 00, and therefore $R_{k}\to 0$.
@@
it implies that the density of $z$ in $E$ is positive, contradicting our assumption that $z$ is a density point of the complementary set ${\mathbb{C}}\setminus E$. Therefore $E$ has full measure. ∎
@@
Because $\Psi^{l}$ is univalent for all $l$, its image is outside a large disk in some ${\mathcal{Z}}^{i}$ plane. Thus, the distortion factor $D$ on each square $S^{\prime}$ of side $\pi$ is close to $1$. Thus we can assume that $ND/(N-1)<1$; for example
@@
Thus $m(\mathcal{W}_{i})=0$, and if $W=\cup_{i=1}^{N}Z_{i}^{-1}(\mathcal{W}_{i})$, then $m(W)=0$; this implies $\cup_{n=1}^{\infty}f^{-n}(W)$ also has zero measure.
To complete the proof, note that $\{z\in E:\omega_{f}(z)\subsetneq P_{f}\}\subset\cup_{n=1}^{\infty}f^{-n}(W)$, so that it has zero measure. ∎
@@
Suppose there is an flimit-$f-$invariant absolutely continuous measure $\rho$.
@@
Moreover, since $\rho$ is also flimit-$f-$invariant, there is an $r>0$ such that $\rho(\mathbf{W}(1))=r$. Furthermore, since $f^{k(p+1)}(\mathbf{W}(1))\subset\mathbf{W}(k)$ for each $k$, the invariance implies that $\rho(\mathbf{W}(k))\geq\rho(\mathbf{W}(1))=r$.

### v3

@@ 3. Disjoint wandering sets of positive Lebesgue measure
Because $\Psi_{i,j}$ is univalent on $G(D_{k})$, the radius of $G(D_{k})$ is less than $\pi/2$; that is,
@@
For each $k>0$, let $\mathcal{G}^{i}_{k}\subset\mathcal{G}^{i}$ be the collection of $S_{m,n}^{i}\subset Hor_{k}^{i}$. It will be convenient to use the index $\delta^{i}$ where $S_{\delta^{i}}=S_{m,n}^{i}$. Note that since $x_{N}^{i}$ is fixed, each such square lies between the vertical lines
$V_{n}=V(x^{i}_{N}+n\pi)=\{{\mathcal{Z}}=x^{i}_{N}+n\pi+iy,y>\alpha_{0}\}$
and
$V_{n+1}=V(x^{i}_{N}+(n+1)\pi)=\{{\mathcal{Z}}=x^{i}_{N}+(n+1)\pi+iy),y>\alpha_{0}\}.$
@@
By lemmas 2.6 and 2.7, for each $j=1,\ldots,N$, $\Psi_{i,j}$ maps the rectangle ${\mathcal{R}}_{\delta^{i}}^{j}$ onto a simply connected domain contained in the horizontal strip $Hor_{k+1}^{j}$.
Let ${\mathcal{U}}^{j}(S_{\delta^{i}})$ denote the union of all the squares $S_{\delta^{j}}$ entirely contained in the interior of $\Psi_{i,j}({\mathcal{R}}_{\delta^{i}}^{j})$:
@@
. Note that because all the squares have side length $\pi$,
@@
Figure 6. The definition of $\mathcal{P}_{\delta^{i}}^{j}$
For each $j=1,\ldots,N$, let ${\mathcal{P}}_{\delta^{i}}^{j}={\mathcal{P}}^{j}_{\delta^{i}}(S_{\delta^{i}})=\Psi_{i,j}^{-1}({\mathcal{U}}^{j}(S_{\delta_{0}^{j}}))$. Since it is the pullback of a union of squares, the following inequality shows that each ${\mathcal{P}}^{j}_{\delta^{i}}$ is a domain whose area is very close to that of the rectangle ${\mathcal{R}}_{\delta^{i}}^{j}$, see Figure 6. Proposition 3.1 shows that the expansion factor for $\Psi_{i,j}$ on $S_{\delta^{i}}$ is at least $\alpha_{k}/(4\pi)$; this says that
@@
Denote the family of indices $({\boldsymbol{\iota}},{\boldsymbol{\sigma}})$ for which ${\mathcal{P}}_{\boldsymbol{\iota}}^{\boldsymbol{\sigma}}$ has a nonempty interior, and is contained in $Hor_{k}^{i_{0}}$, by $\mathcal{I}_{k}^{l}$ . For readability, identify the index $(\boldsymbol{\iota},{\boldsymbol{\sigma}})\in\mathcal{I}_{k}^{l}$ with the set of points in ${\mathcal{P}}_{\boldsymbol{\iota}}^{\boldsymbol{\sigma}}$.
@@
$Y^{l}=\cup_{(\boldsymbol{\iota},\boldsymbol{\sigma})\in\mathcal{I}^{l}}{\mathcal{P}}_{\boldsymbol{\iota}}^{\boldsymbol{\sigma}}.$
Note that, by definition, $Y^{l+1}\subset Y^{l}$. Let
(12) $Y^{\infty}=\cap_{l=0}^{\infty}Y^{l}.$
@@
By definition, if $K$ is the set defined by $\mathcal{I}^{l}_{k}$, then $\Psi^{l+1}(K)$ is a union of squares in $\cup_{\delta^{j}\in\Delta}S_{\delta^{j}}$ in $Hor_{k+1+l}^{j_{l}}$. Denote this union by $W=\Psi^{l+1}(K)$. Recall that each square $S_{\delta^{j}}$ contains the sets $\mathcal{P}_{\delta^{j}}^{i},i=1,\ldots,N$, each of which is mapped to a square, Therefore,
$\Psi^{l+1}(W\cap Y^{l+1})=\cup_{j_{l}\in\Delta}\cup_{i=1}^{N}{\mathcal{P}}_{S_{\delta^{j_{l}}}}^{i}.$
The union of the orbits of the asymptotic values $\lambda_{i}$ is a finite set. Since the points ${\mathcal{Z}}_{j_{1}}^{-1}(Hor_{k+1+l}^{j_{l}})$ lie in the asymptotic tract of $\lambda_{i}$, we can choose $k$ so that they are far from orbit points. In fact, we can choose it so that there is a neighborhood of ${\mathcal{Z}}_{j_{1}}^{-1}(Hor_{k+1+l}^{j_{l}})$ of radius $2\pi$ disjoint from the orbit points. It follows that the branches of $\Psi^{-(l+1)}$ are well defined on this neighborhood.
Denote the branch that maps $W$ back to $K$ by $\Xi$ and apply the Koebe distortion theorem. It says that for $\zeta,\xi\in W$, there exists a constant $C_{0}$ independent of $\Xi$, such that
@@
$\frac{m(W\setminus Y^{l+1})}{m(W)}=\frac{m(\Xi(\cup(S_{\delta^{j_{l}}}\setminus\cup_{i=1}^{N}{\mathcal{P}}_{\delta_{j_{l}}}^{i})))}{m(\Xi(\cup S_{\delta^{j_{l}}}))}\leq C_{0}^{2}\frac{\sum m(S_{\delta_{j_{l}}}\setminus\cup_{i=1}^{N}{\mathcal{P}}_{\delta^{j_{l}}}^{i})}{\sum m((S_{\delta^{j_{1}}})}\leq\frac{C}{\alpha_{|k|+l+1}},$
for some constant $C=(4\sqrt{2}\pi^{2}+6)C_{0}^{2}>0$. Note that the last inequality follows from Equation(11). ∎
@@
    Every pair of points $({\mathcal{Z}},{\mathcal{Z}}^{\prime})\in(S_{\delta^{i_{0}}},S_{\delta^{i_{0}}}^{\prime}),$ satisfies $|\Im{\mathcal{Z}}|-|\Im{\mathcal{Z}}^{\prime}|\geq 2\alpha_{k-1}$.
@@
    $\Psi^{l}(W_{1})\subset Hor^{i_{1}}_{k+l}$, $\Psi^{l}(W_{2})\subset Hor^{i_{2}}_{k+l}$ for some $i_{1},i_{2}\in\{1,\ldots,N\}$;
@@
The first assertion shows that both $W_{1}$ and $W_{2}$ are wandering sets and the second shows that their forward orbits are disjoint. Note that ${\mathcal{Z}}_{i_{0}}$ is analytic, so both ${\mathcal{Z}}_{i_{0}}^{-1}(W_{1}),{\mathcal{Z}}_{i_{0}}^{-1}(W_{2})$ have positive Lebesgue measure.
@@ 4. Main Theorems 2 and 3
Let $P_{f}=\cup_{i=1}^{N}\{\lambda_{i},f(\lambda_{i}),\ldots,f^{p_{i}-1}(\lambda_{i}),\infty\}$ be the post-singular set. As we proved above, $f$ is not ergodic. From the extended dichotomy of Bock [Boc], we know that for almost every point $z\in{\mathbb{C}}$, $\lim_{n\to\infty}d(f^{n}(z),P_{f})\to 0$. This implies that the $\omega$-limit set $\omega(z)\subset P_{f}$ and in terms of the auxiliary variables, it says that for each such $z$, there exists a fixed $i\in\{1,\ldots,N\}$, a sequence $n_{k}\to\infty$, and an $n_{k_{0}}$ such that for all $n_{k}\geq n_{k_{0}}$, $f^{n_{k}}(z)$ is in the asymptotic tract $T_{i}$ and $\Im{\mathcal{Z}}_{i}(f^{n_{k}}(z))\to\infty$.
The set $Y^{\infty}$ was defined in (12) as a subset of $\cup_{i=1}^{N}{\mathcal{Z}}_{i}$ and we showed in lemma 3.3 that it has positive Lebesgue measure. In order to prove Main Theorem 2, which is a result about points in $\mathbb{C}$}, we need to pull $Y^{\infty}$ back to $\mathbb{C}$}. To do this, define $E^{\infty}=\cup_{i=1}^{N}{\mathcal{Z}}_{i}^{-1}(Y^{\infty}\cap{\mathcal{Z}}_{i})$ and set $E=\cup_{n\in{\mathbb{N}}}f^{-n}(E^{\infty})$.
@@
The set $E$ in $\mathbb{C}$} has full Lebesgue measure; that is, $m({\mathbb{C}}\setminus E)=0$.
@@
Note that these inequalities define a set of strips $\widetilde{Hor^{i}_{l_{k}}}$ that is different from, but intersects the strips $Hor^{i}_{n_{k}}$.
@@
Because the diameter of $\widetilde{Q}^{n_{k}}$ is less than $16\sqrt{2}\alpha_{k-1}$, the diameter of $F_{n_{k}}(\widetilde{Q}^{n_{k}})$ tends to 0, and therefore $R_{k}\to 0$.
@@
it implies that the density of $z$ in $E$ is positive, contradicting our assumption that $z$ is a density point of the complementary set ${\mathbb{C}}\setminus E$. Therefore $E$ has full Lebesgue measure. ∎
@@
Because $\Psi^{l}$ is univalent for all $l$, its image is outside a large disk in some ${\mathcal{Z}}^{i}$ plane. Thus, the distortion factor $D$ on each square $S^{\prime}$ of side $\pi$ is close to $1$. Thus we can assume that ND/(N1)<1ND/(N-1)<1; for example
@@
Thus $m(\mathcal{W}_{i})=0$, and if $W=\cup_{i=1}^{N}Z_{i}^{-1}(\mathcal{W}_{i})$, then $m(W)=0$; this implies $\cup_{n=1}^{\infty}f^{-n}(W)$ also has zero Lebesgue measure.
To complete the proof, note that $\{z\in E:\omega_{f}(z)\subsetneq P_{f}\}\subset\cup_{n=1}^{\infty}f^{-n}(W)$, so that it has zero Lebesgue measure. ∎
@@
Suppose there is an $f$-invariant absolutely continuous measure $\rho$.
@@
Moreover, since $\rho$ is also $f$-invariant, there is an $r>0$ such that $\rho(\mathbf{W}(1))=r$. Furthermore, since $f^{k(p+1)}(\mathbf{W}(1))\subset\mathbf{W}(k)$ for each $k$, the invariance implies that $\rho(\mathbf{W}(k))\geq\rho(\mathbf{W}(1))=r$.